OpenStudy (anonymous):

how do you find the measure of an angle that is four hundred twenty less than three times its supplement?

hero (hero):

(3x - 420) + x = 180

hero (hero):

4x = 600

hero (hero):

x = 150

hero (hero):

other angle is 30

hero (hero):

I was going at such a snail's pace. Fiddle you could have easily gotten this one if you wanted it.

OpenStudy (anonymous):

I did want to. I really did ....

hero (hero):

lol

OpenStudy (anonymous):

still a bit confused about this -420 degree difference, since it's cyclic and all.

hero (hero):

I think you're thinking into it way too much. Think algebra for these. We're only talking about two angles that add up to 180

OpenStudy (anonymous):

you guys are so smart

hero (hero):

I'm not smart at all

hero (hero):

Fiddle might be smarter than me....

OpenStudy (anonymous):

lol you answred it first

hero (hero):

Who answers first isn't necessary a measure of smartness

hero (hero):

I got here first

OpenStudy (anonymous):

don't we need to do 420-360 = 60 or something ?

hero (hero):

Why would we need to do that? You're thinking about trig and not geometry. You have to know when to separate the two.

OpenStudy (anonymous):

maybe not...

OpenStudy (anonymous):

you're right - you're smarter.

hero (hero):

One question isn't necessarily a measure of smartness

OpenStudy (anonymous):

I was overcomplicating it. :)

OpenStudy (anonymous):

Well its not a race anyways :)

OpenStudy (anonymous):

so its 3x +x =4x and then wat

OpenStudy (anonymous):

okay so dont answer me -.-

OpenStudy (anonymous):

Hero found the angle was 150

hero (hero):

I thought I posted the solution already

OpenStudy (anonymous):

yeah but i wanna understand how you did it. you know

OpenStudy (anonymous):

im so confused here. i was thinking you do 3x-420+x=180

hero (hero):

Check the first line of this board....

OpenStudy (anonymous):

okay. i did i juss dont understand whwere you got 4x=600 ?

hero (hero):

3x + x = 4x 180 + 420 = 600

OpenStudy (anonymous):

ohhhhh

hero (hero):

:o)